OpenStudy (anonymous):

what is log8 ove loge 9

OpenStudy (jack1):

hi @lasya

OpenStudy (jack1):

assuming its log (so base 10) not natural log (ln) \[ \Large\frac {\log_{}8}{ \log_{}9} \] \[= \Large\frac {\log_{}2^3}{ \log_{}3^2} \] \[= \Large\frac {3 \log_{}2}{ 2 \log_{}3} \] \[= \Large ~about ~0.9464 \]

OpenStudy (irishboy123):

\(\large log_9 8 \)